I Need it by Hand and not with Excel since I have no idea About Excel and cannot use it on the exam. Thanks in advance

Epple Computer Corp. (ECC) went public 8 years ago and issued 15 Million shares at a nominal value
€ 10 per share. At present, there is a need for additional funding to further grow the company. According
to the business plan, the expected return on the total capital invested in ECC is averaging at 9 %. The
company could issue another bond at 8,5 % interest rate or issue new shares.
a. The management of ECC suggests an increase of share capital at the Shareholders Meeting
with the following specifications: (6 points)
 Additional number of shares to be issued: 3 Million
 The current share price (before share capital increase) is € 110 per share.
 The new shares shall be issued in the stock market at a price of € 95 per share.
 The shares shall be issued with subscription rights offered to the current
shareholders.
Please calculate the share price after the share capital increase and the value of a
subscription right.
Answer #1

stock price after issued = (current market cap+amount raised)/(current shares+additional shares)

=(110*15+95*3)/(15+3) = 107.5

Value of rights = (current price-subscription price)/( number of rights required to purchase one share) = (110-95)/(15/3) = 3